darryl144:

If the APR of a savings account is 6.4% and interest is compounded quarterly, what is the approximate APY of the account?

darryl144:

A. 6% B. 6.56% C. 1.05% D. 10.49%

Elsa213:

`You should apply the formula,` `APY=100((1+APR/Periods)Periods−1)` `In this case,` `APY=100((1+0.064/4)4−1)=6.56%` Source: questioncove.com/updates/5237a868e4b0b9d6b2d6c109